The number of customers arriving per hour at a certain automobile service facility is assumed to follow a Poisson distribution with mean λ=7.

​(a) Compute the probability that more than 9 customers will arrive in a 2​-hour period.

​(b) What is the mean number of arrivals during a 2​-hour ​period?


1
Expert's answer
2021-10-28T07:02:14-0400

Let "X=" the number of customers arriving per period "t" : "X\sim Po(\lambda t)."

(a) Given "\lambda=7, t=2\ h"


"\lambda t=7(2)=14"


"P(X>9)=1-P(X\leq9)"

"=1-P(X=0)-P(X=1)-P(X=2)"

"-P(X=3)-P(X=4)-P(X=5)"

"-P(X=6)-P(X=7)-P(X=8)"

"-P(X=9)=1-\dfrac{e^{-14}\cdot14^0}{0!}-\dfrac{e^{-14}\cdot14^1}{1!}"

"-\dfrac{e^{-14}\cdot14^2}{2!}-\dfrac{e^{-14}\cdot14^3}{3!}-\dfrac{e^{-14}\cdot14^4}{4!}"

"-\dfrac{e^{-14}\cdot14^5}{5!}-\dfrac{e^{-14}\cdot14^6}{6!}-\dfrac{e^{-14}\cdot14^7}{7!}"

"-\dfrac{e^{-14}\cdot14^8}{8!}-\dfrac{e^{-14}\cdot14^9}{9!}=0.89060063"

(b)


"mean=\lambda t=7(2)=14"

The mean number of arrivals during a 2​-hour ​period is "14."